# MagicPose4D: Crafting Articulated Models with Appearance and Motion Control

### v2

@@ 1 Introduction
Dual-Phase 4D Reconstruction Module: Given the complexity, especially with articulated objects, we divide this module into two phases. First, we use image descriptors (e.g., segments, flow) as 2D supervision. Additionally, we employ relatively less accurate but geometrically informative pseudo-3D models as geometric priors (using image-to-3D models) to provide 3D supervision for 3D reconstruction. This phase does not impose constraints emanating from the articulated nature of the object, e.g., captured in the underlying skeletal structure and plausible changes in it during motion, allowing each part to learn arbitrary rotations ($\mathbf{R}$) and translations ($\mathbf{t}$) while focusing on learning the model’s shape. In the second phase, we ensure the physical plausibility of the motion by enforcing kinematic chain-based skeletal movement constraints. Additionally, we propose a Global-local Chamfer loss, which ensures that the overall distribution of predicted mesh vertices aligns with the supervision while maintaining part-level alignment, without requiring additional annotations. Cross-category Motion Transfer Module: This module achieves cross-species motion transfer by mapping the skeleton of one species to another by establishing joint and limb correspondences through a kinematic-chain-based representation of the skeleton. Our motion transfer module is non-training-based. It helps improve generalization and prevents the poor performance seen in many existing methods when tested on data with significant gaps from the training set. Also, we leverage dynamic rigidity Zhang et al. (2024c) to guarantee smoothness between frames, unlike the existing approaches Song et al. (2021; 2023), which perform frame-independent pose transfer.
@@ 2 Related Work & Motivation
4D/3D Reconstruction. Significant advancements in 4D/3D reconstruction include the development of specialized parametric models such as SCAPE Allen et al. (2003) and SMPL Loper et al. (2023a) for human bodies, MANO Romero et al. (2022) for hand movements, FLAME Li et al. (2017) and EMOCA Daněček et al. (2022) for facial expressions, and SMAL Zuffi et al. (2017) for quadruped animals. However, these methods require predefined parameters that include skeletons and skinning weights, limiting their generalization to uncommon species and out-of-distribution data. Recent neural implicit-based methods Palafox et al. (2021); Yang et al. (2022b); Giebenhain et al. (2023); Zhang et al. (2024c); Yang et al. (2021b) offer promising alternatives by jointly learning static 3D meshes and time-varying parameters without predefined templates. However, they often fail with monocular videos containing sparse views, limiting their practical applicability. To address this, MagicPose4D introduces an innovative Dual-Phase 4D Reconstruction module, which reduces the requirements from multi-view videos to single-view videos and achieves robust reconstruction.
Diffusion-based 4D Generation and Pose Transfer. Recent diffusion-based 4D generation methods have shown promising results by leveraging text-to-3D models followed by text-to-video supervision. These techniques Singer et al. (2023); Ren et al. (2023); Ling et al. (2023); Bahmani et al. (2023); Zheng et al. (2023); Gao et al. (2024); Yang et al. (2024); Jiang et al. (2023); Zhao et al. (2023) have improved the geometry and appearance of generated models but are generally limited to minor movements within a fixed location and rely on text prompts, making precise motion control difficult. Our goal is to create dynamic 4D animations that closely transfer the motion of any given reference real-world object. For pose transfer, recent skeleton-based frameworks Song et al. (2021); Liao et al. (2022) have explored the use of rigging points and key points. However, learning-based methods often struggle with in-the-wild data and significant domain gaps between identities. MagicPose4D introduces a cross-category motion transfer module that supports cross-species transfer while ensuring generalization and maintaining temporal smoothness.
@@ 3.2.2 Supervision and Losses
As shown in Fig. 2, given the input from a monocular video, we utilize segmentation models Kirillov et al. (2023); Zhang et al. (2024a) and optical flow prediction models Teed & Deng (2020) to obtain 2D supervision. Furthermore, we employ an image-to-3D model to independently predict meshes for each frame, serving as pseudo-3D supervision. A question arises: "Why not directly use an image-to-3D model to independently predict 3D meshes for each frame?" As illustrated in Fig.6 (c), objects always exhibit self-occlusion in the video, under which circumstances image-to-3D models typically fail to produce accurate results. For instance, some frames might only depict a camel with two visible legs, resulting in a 3D mesh sequence that does not effectively capture the action information portrayed in the video. Moreover, as demonstrated in Video 3 of the supplementary materials, meshes generated directly from the image-to-3D module are independent of one another, thus lacking temporal continuity and smoothness. Although the initial pseudo-3D supervision may not be very accurate, it still provides valuable geometric priors that help address the information loss caused by insufficient perspectives of the object in the video.
@@ 3.3 Cross-Category Motion Transfer
In the second phase of 4D Reconstruction, as shown in Fig. 3(a), we extract the underlying skeletal motion of the reference meshes. Given a sequence of meshes for the reference object, the skeletonization module extracts the skeleton of the canonical shape, which is defined as the shape corresponding to time $=0$. We fix the canonical skeleton and the skinning weights, thus controlling the skeleton’s pose solely by learning the pre-frame angles and bone scales. Subsequently, by employing blend skinning, we obtain the deformed mesh.
The input to the skeletonization module is a mesh, along with an optional skeleton template. When the reference object is a commonly recognized form, such as a quadruped or a human, we utilize the corresponding skeleton template and embed it into the mesh. Following the methodology described in Baran & Popović (2007), we construct a coarse discrete representation to locate the approximate position of the skeleton within the internal space of the mesh. This process involves embedding the skeleton into a graph derived from the character’s internal volume and determining an optimal solution using the A* algorithm across all possible matches. When a template of the object is not available, we employ skeleton extraction methods Bærentzen & Rotenberg (2021); Zhang et al. (2024c) to extract the canonical skeleton of the reference object. After learning the reference motion represented by the pre-frame angles and the bone scale of the skeleton, we can readily compute the global-local transformation of the target object using forward kinematics, as illustrated in Fig. 3(b). We extract the canonical skeleton for the target object by inputting the canonical skeleton of the reference object, adhering to the method previously described following Baran & Popović (2007). Subsequently, blend skinning is employed to generate the deformed meshes.
@@ 4 Experiments
Since the primary goal of our 4D reconstruction module is to better capture motion, we focus on the accuracy of keypoint reconstruction. Therefore, we use keypoint transfer accuracy as the quantitative evaluation metric. We examine the performance of MagicPose4D against the SOTA methods, S3O Zhang et al. (2024b), LIMR Zhang et al. (2024c), LASR Yang et al. (2021a) and ViSER Yang et al. (2021c). As shown in Tab.2, MagicPose4D consistently outperforms them for all animal subjects with a notable margin. When we run the author-provided codes for LASR, and ViSER on DAVIS, and PlanetZoo, we find that their results are highly variable across different runs. The results we report for the mean accuracies we have obtained over multiple runs ($5$) of all methods. Furthermore, as depicted in Fig.7, we compare the results of MagicPose4D with LASR Yang et al. (2021a) and BANMo Yang et al. (2022a). These methods fall short of achieving ideal reconstruction, due to a lack of structural information about the objects, since they learn from only a single monocular video, which contains sparse views. NeRF-based methods such as BANMo Yang et al. (2022b) and MagicPony Wu et al. (2023) require particularly dense views from multiple cameras or long videos to achieve decent results, while ours achieves good reconstructions from only sparse views.
@@
Table 2: 2D Keypoint Transfer Accuracy. This table presents the 2D keypoint transfer accuracy on DAVIS and PlanetZoo datasets. For all methods, we carry out multiple executions and report the mean of the observed accuracy in each case. Note that all experiment results of MP4D are without the skeleton templates and learning underlying structure following Zhang et al. (2024b).
@@
LIMR Zhang et al. (2024c) 77.5 79.6 65.4 84.7 79.6 59.3 64.9 68.3 64.1 57.1 55.2 68.7
S3O Zhang et al. (2024b) 79.8 83.2 67.3 86.7 84.7 63.1 72.8 71.3 66.4 61.5 62.1 72.6
@@ 5 Conclusion & Limitations
The main limitations of our method and existing works are: (i) Deformation based on blend skinning relies on accurate and robust skeletons and skinning weights predictions, facing a trade-off between generalization and accuracy. Learning-based methods have limited generalization due to restricted training datasets, while non-learning methods suffer from inductive bias, leading to suboptimal results. (ii) MagicPose4D can infer poses quickly for pose transfer without training, but 4D reconstruction requires significant training (5 hours on an L40S). (iii) Our method struggles with detailed motion control, such as fingers and facial features, due to the challenge of capturing fine-grain details during 4D reconstruction. These issues represent future research directions.
@@ A.5 User Study
Similarly, we provide a user study for comparison with 4Dfy, DMT+Stag4D, and SVD+Stag4D. In each comparison experiment, we collect reference images, text descriptions of motion, mesh reference motion sequences, and generated videos from both methods. For 4Dfy we follow the official code and instructions, for the other two, we first use SVD to generate a video via text prompt or use DMT to do video-level motion transfer and then use Stag4D to do reconstruction from the generated video. For MagicPose4D, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence. We compared the rendered videos from Animate124 to videos of generated mesh from MagicPose4D in different viewpoints. Criteria for judgment: 1) The generation’s identity should match the reference image. 2) The generated mesh sequence should be consistent and smooth. From the results presented in Tab. 1, we conclude that MagicPose4D provides the most satisfying visualizations.
@@
The 3D loss is described in the main article. Here we introduce the 2D losses and regularization terms. The 2D losses are similar to those in existing differentiable rendering pipelines Yang et al. (2021b; 2022b); Zhang et al. (2024c). We define $\mathbf{S^{t}}$,$\mathbf{I^{t}}$, $\mathbf{F}^{2\text{D},t}$ as the silhouette, input image, and optical flow of the input image, and their corresponding rendered counterparts as $\mathbf{\tilde{S}^{t}}$,$\mathbf{\tilde{I}^{t}}$, $\mathbf{\tilde{F}}^{2\text{D},t}$. The following losses ensure the fitting between rendered and original:
@@
where pdist(,) is the perceptual distance Zhang et al. (2018). Also, we leverage three regularization terms: (1) Dynamic Rigidity term, which is introduced in paper.Zhang et al. (2024c). (2) Symmetry term, which encourages the canonical mesh to be symmetric.

### v3

@@ 1 Introduction
Dual-Phase 4D Reconstruction Module: Given the complexity, especially with articulated objects, we divide this module into two phases. First, we use image descriptors (e.g., segments, flow) as 2D supervision. Additionally, we employ relatively less accurate but geometrically informative pseudo-3D models as geometric priors (using image-to-3D models) to provide 3D supervision for 3D reconstruction. This phase does not impose constraints emanating from the articulated nature of the object, e.g., captured in the underlying skeletal structure and plausible changes in it during motion, allowing each part to learn arbitrary rotations ($\mathbf{R}$) and translations ($\mathbf{t}$) while focusing on learning the model’s shape. In the second phase, we ensure the physical plausibility of the motion by enforcing kinematic chain-based skeletal movement constraints. Additionally, we propose a Global-local Chamfer loss, which ensures that the overall distribution of predicted mesh vertices aligns with the supervision while maintaining part-level alignment, without requiring additional annotations. Cross-category Motion Transfer Module: This module achieves cross-species motion transfer by mapping the skeleton of one species to another by establishing joint and limb correspondences through a kinematic-chain-based representation of the skeleton. Our motion transfer module is non-training-based. It helps improve generalization and prevents the poor performance seen in many existing methods when tested on data with significant gaps from the training set. Also, we leverage dynamic rigidity Zhang et al. (2024d) to guarantee smoothness between frames, unlike the existing approaches Song et al. (2021; 2023), which perform frame-independent pose transfer.
@@ 2 Related Work & Motivation
4D/3D Reconstruction. Significant advancements in 4D/3D reconstruction include the development of specialized parametric models such as SCAPE Allen et al. (2003) and SMPL Loper et al. (2023a) for human bodies, MANO Romero et al. (2022) for hand movements, FLAME Li et al. (2017) and EMOCA Daněček et al. (2022) for facial expressions, and SMAL Zuffi et al. (2017) for quadruped animals. However, these methods require predefined parameters that include skeletons and skinning weights, limiting their generalization to uncommon species and out-of-distribution data. Recent neural implicit-based methods Palafox et al. (2021); Yang et al. (2022b); Giebenhain et al. (2023); Zhang et al. (2024d); Yang et al. (2021b) offer promising alternatives by jointly learning static 3D meshes and time-varying parameters without predefined templates. However, they often fail with monocular videos containing sparse views, limiting their practical applicability. To address this, MagicPose4D introduces an innovative Dual-Phase 4D Reconstruction module, which reduces the requirements from multi-view videos to single-view videos and achieves robust reconstruction.
Diffusion-based 4D Generation and Pose Transfer. Recent diffusion-based 4D generation methods have shown promising results by leveraging text-to-3D models followed by text-to-video supervision. These techniques Singer et al. (2023); Ren et al. (2023); Ling et al. (2023); Bahmani et al. (2023); Zheng et al. (2023); Gao et al. (2024); Yang et al. (2024); Jiang et al. (2023); Zhao et al. (2023); Chen et al. (2024; 2023) have improved the geometry and appearance of generated models but are generally limited to minor movements within a fixed location and rely on text prompts, making precise motion control difficult. Our goal is to create dynamic 4D animations that closely transfer the motion of any given reference real-world object. For pose transfer, recent skeleton-based frameworks Song et al. (2021); Liao et al. (2022) have explored the use of rigging points and key points. However, learning-based methods often struggle with in-the-wild data and significant domain gaps between identities. MagicPose4D introduces a cross-category motion transfer module that supports cross-species transfer while ensuring generalization and maintaining temporal smoothness.
@@ 3.2.2 Supervision and Losses
As shown in Fig. 2, given the input from a monocular video, we utilize segmentation models Kirillov et al. (2023); Zhang et al. (2024b) and optical flow prediction models Teed & Deng (2020) to obtain 2D supervision. Furthermore, we employ an image-to-3D model to independently predict meshes for each frame, serving as pseudo-3D supervision. A question arises: "Why not directly use an image-to-3D model to independently predict 3D meshes for each frame?" As illustrated in Fig.6 (c), objects always exhibit self-occlusion in the video, under which circumstances image-to-3D models typically fail to produce accurate results. For instance, some frames might only depict a camel with two visible legs, resulting in a 3D mesh sequence that does not effectively capture the action information portrayed in the video. Moreover, as demonstrated in Video 3 of the supplementary materials, meshes generated directly from the image-to-3D module are independent of one another, thus lacking temporal continuity and smoothness. Although the initial pseudo-3D supervision may not be very accurate, it still provides valuable geometric priors that help address the information loss caused by insufficient perspectives of the object in the video.
@@ 3.3 Cross-Category Motion Transfer
In the second phase of 4D Reconstruction described in Sec. 3.2, as shown in Fig. 3(a), we extract the underlying skeletal motion of the reference meshes. Given a sequence of meshes for the reference object, the skeletonization module extracts the skeleton of the canonical shape, which is defined as the shape corresponding to time $=0$. We fix the canonical skeleton and the skinning weights, thus controlling the skeleton’s pose solely by learning the pre-frame angles and bone scales (Sec.3.2.1). Subsequently, by employing blend skinning, we obtain the deformed mesh for the reference object.
The input to the skeletonization module is a mesh, along with an optional skeleton template. When the reference object is a commonly recognized form, such as a quadruped or a human, we utilize the corresponding skeleton template and embed it into the mesh. Following the methodology described in Baran & Popović (2007), we construct a coarse discrete representation to locate the approximate position of the skeleton within the internal space of the mesh. This process involves embedding the skeleton into a graph derived from the character’s internal volume and determining an optimal solution using the A* algorithm across all possible matches. When a template of the object is not available, we employ skeleton extraction methods Bærentzen & Rotenberg (2021); Zhang et al. (2024d) to extract the canonical skeleton of the reference object. After learning the reference motion represented by the pre-frame angles and the bone scale of the skeleton from the reconstruction module, we can readily compute the global-local transformation of the target object using forward kinematics, as illustrated in Fig. 3(b). We extract the canonical skeleton for the target object by inputting the canonical skeleton of the reference object, adhering to the method previously described following Baran & Popović (2007). Subsequently, blend skinning is employed to generate the deformed meshes.
@@ 4 Experiments
Since the primary goal of our 4D reconstruction module is to better capture motion, we focus on the accuracy of keypoint reconstruction. Therefore, we use keypoint transfer accuracy as the quantitative evaluation metric. We examine the performance of MagicPose4D against the SOTA methods, S3O Zhang et al. (2024c), LIMR Zhang et al. (2024d), LASR Yang et al. (2021a) and ViSER Yang et al. (2021c). As shown in Tab.2, MagicPose4D consistently outperforms them for all animal subjects with a notable margin. When we run the author-provided codes for LASR, and ViSER on DAVIS, and PlanetZoo, we find that their results are highly variable across different runs. The results we report for the mean accuracies we have obtained over multiple runs ($5$) of all methods. Furthermore, as depicted in Fig.7, we compare the results of MagicPose4D with LASR Yang et al. (2021a) and BANMo Yang et al. (2022a). These methods fall short of achieving ideal reconstruction, due to a lack of structural information about the objects, since they learn from only a single monocular video, which contains sparse views. NeRF-based methods such as BANMo Yang et al. (2022b) and MagicPony Wu et al. (2023) require particularly dense views from multiple cameras or long videos to achieve decent results, while ours achieves good reconstructions from only sparse views.
@@
Table 2: 2D Keypoint Transfer Accuracy. This table presents the 2D keypoint transfer accuracy on DAVIS and PlanetZoo datasets. For all methods, we carry out multiple executions and report the mean of the observed accuracy in each case. Note that all experiment results of MP4D are without the skeleton templates and learning underlying structure following Zhang et al. (2024c).
@@
LIMR Zhang et al. (2024d) 77.5 79.6 65.4 84.7 79.6 59.3 64.9 68.3 64.1 57.1 55.2 68.7
S3O Zhang et al. (2024c) 79.8 83.2 67.3 86.7 84.7 63.1 72.8 71.3 66.4 61.5 62.1 72.6
@@ 5 Conclusion & Limitations
The main limitations of our method and existing works are: (i) Deformation based on blend skinning relies on accurate and robust skeletons and skinning weights predictions, facing a trade-off between generalization and accuracy. Learning-based methods have limited generalization due to restricted training datasets, while non-learning methods suffer from inductive bias, leading to suboptimal results. (ii) MagicPose4D can infer poses quickly for pose transfer without training, but 4D reconstruction requires significant training (5 hours on an L40S). (iii) Since our method performs explicit pose transfer between two objects with similar skeletal structures, the resulting motion may appear unnatural when the target object differs significantly from the reference object in terms of skeleton or shape. This limitation highlights a promising direction for future research: developing motion transfer techniques that operate across multiple levels of abstraction. Instead of relying solely on direct pose-level transfer, it would be valuable to explore approaches that can perform transfer at higher levels (e.g., activity-level) or intermediate levels (e.g., primitive activity-level).
@@ A.5 User Study
Similarly, we provide a user study for comparison with 4Dfy, DMT+Stag4D, and SVD+Stag4D. In each comparison experiment, we collect reference images, text descriptions of motion, mesh reference motion sequences, and generated videos from both methods. For 4Dfy we follow the official code and instructions, for the other two, we first use SVD to generate a video via text prompt or use DMT to do video-level motion transfer and then use Stag4D to do reconstruction from the generated video. For MagicPose4D, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence. We compared the rendered videos from Animate124 to videos of generated mesh from MagicPose4D in different viewpoints. Criteria for judgment: 1) The generation’s identity should match the reference image. 2) The generated mesh sequence should be consistent and smooth. From the results presented in Tab. 1, we conclude that MagicPose4D provides the most satisfying visualizations. Zhang & Fu (2025); Zhang et al. (2024a; 2023b; 2021b; 2023a; 2021c; 2021a)
@@
The 3D loss is described in the main article. Here we introduce the 2D losses and regularization terms. The 2D losses are similar to those in existing differentiable rendering pipelines Yang et al. (2021b; 2022b); Zhang et al. (2024d). We define $\mathbf{S^{t}}$,$\mathbf{I^{t}}$, $\mathbf{F}^{2\text{D},t}$ as the silhouette, input image, and optical flow of the input image, and their corresponding rendered counterparts as $\mathbf{\tilde{S}^{t}}$,$\mathbf{\tilde{I}^{t}}$, $\mathbf{\tilde{F}}^{2\text{D},t}$. The following losses ensure the fitting between rendered and original:
@@
where pdist(,) is the perceptual distance Zhang et al. (2018). Also, we leverage three regularization terms: (1) Dynamic Rigidity term, which is introduced in paper.Zhang et al. (2024d). (2) Symmetry term, which encourages the canonical mesh to be symmetric.